# An Analysis under a Unified Formulation of Learning Algorithms with Output Constraints

### v1

@@ Abstract
[…]
@@ 3 Further exploration on integration of main task and constraint information
, whenever $\nabla\mathcal{C}(\theta)\cdot g_{sup}<0$. Then, the vector $\text{Proj}(\nabla\mathcal{C}(\theta))$ satisfies $\text{Proj}(\nabla\mathcal{C}(\theta))\cdot g_{sup}=0$. This ensures that $\nabla\mathcal{C}$ is transformed orthogonally to $g_{sup}$, preventing ti from providing information that contradicts supervised learning.
@@
, whenever $\nabla\mathcal{L}(\theta)\cdot g_{con}<0$. Then, the vector $\text{Proj}(\nabla\mathcal{L}(\theta))$ satisfies $\text{Proj}(\nabla\mathcal{L}(\theta))\cdot g_{con}=0$. This ensures that $\nabla\mathcal{L}$ is transformed orthogonally to $g_{con}$, preventing ti from providing information that contradicts constraint injection.
@@ A.1 Fine-grained expressiveness of constraints
Table 1: Examples of logics. Our experiment used Gödel logic except for the implication ($\Rightarrow$). For $\Rightarrow$, we used S-implication form, $\max(1-a,b)$.

### v2

@@ Abstract
Neural networks (NN) excel in diverse tasks but can produce nonsensical results due to their exclusive reliance on (input, output) pairs, often conflicting with human knowledge. Injecting human knowledge via output constraints can enhance performance and reduce violations. Despite attempts to compare existing algorithms, no unified categorization of learning algorithms with output constraints exists. Our contributions are: (1) We categorize previous studies using three axes: type of constraint loss (e.g., probabilistic soft logic, REINFORCE), exploration strategy of constraint-violating examples, and integration mechanism for balancing the main task and constraints learning signals. (2) We propose new algorithms inspired by continual-learning for integrating main task and constraint information. (3) We introduce the $H\beta$-score metric to simultaneously evaluate main task performance and constraint violation. Our experiments on NLP tasks (NLI, STE, SRL) show that our projection-based integration mechanism outperforms others. Sampling strategy is crucial for high $H\beta$-scores, with better results as sample numbers increase. Additionally, soft-type constraint loss performs well when combined with sampling strategies. These insights highlight key factors for achieving high $H\beta$-scores and demonstrate the efficacy of our methods.
@@ 3 Further exploration on integration of main task and constraint information
, whenever $\nabla\mathcal{C}(\theta)\cdot g_{sup}<0$. Then, the vector $\text{Proj}(\nabla\mathcal{C}(\theta))$ satisfies $\text{Proj}(\nabla\mathcal{C}(\theta))\cdot g_{sup}=0$. This ensures that $\nabla\mathcal{C}$ is transformed orthogonally to $g_{sup}$, preventing it from providing information that contradicts supervised learning.
@@
, whenever $\nabla\mathcal{L}(\theta)\cdot g_{con}<0$. Then, the vector $\text{Proj}(\nabla\mathcal{L}(\theta))$ satisfies $\text{Proj}(\nabla\mathcal{L}(\theta))\cdot g_{con}=0$. This ensures that $\nabla\mathcal{L}$ is transformed orthogonally to $g_{con}$, preventing it from providing information that contradicts constraint injection.
@@ 7 Conclusions
8 Limitations and future work
Our experiments were exclusively conducted on NLP tasks (NLI, STE, SRL) and did not include cutting-edge large language models. Therefore, it would be worthwhile to extend our experiments to a broader range of tasks and larger models. This would not only validate the generalizability of our methods but also potentially uncover new insights and improvements for various applications.
Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant (RS-2023-00280883, RS-2023-00222663), by the National Super computing Center with super computing resources including technical support (KSC-2023-CRE-0176), and partially supported by New Faculty Startup Fund from Seoul National University.
@@ A.1 Fine-grained expressiveness of constraints
Table 1: Examples of logics. Our experiment used Gödel logic except for the implication ($\Rightarrow$). For $\Rightarrow$, we used S-implication Baczyński and Jayaram (2007); Bedregal et al. (2010) form, $\max(1-a,b)$.